Bonita Company is considering a long-term investment project called ZIP. ZIP will require an investment of $121,600. It will have a useful life of 4 years and no salvage value. Annual revenues would increase by $78,904, and annual expenses (excluding depreciation) would increase by $40,600. Bonita uses the straight-line method to compute depreciation expense. The company’s required rate of return is 10%.

Compute the annual rate of return.

Annual rate of return

%


Determine whether the project is acceptable?

AcceptReject

the project.
Answer #1

Cost of project = $121,600

Useful life = 4 years

Annual depreciation = Cost of project / Useful life

= 121,600/4

= $30,400

Increase in annual revenue = $78,904

Increase in annual expense = $40,600

Average annual income = Increase in annual revenue - Increase in annual expense - Annual depreciation

= 78,904-40,600-30,400

= $7,904

Average investment = Initial investment/2

= 121,600/2

= $60,800

Annual rate of return = Average annual income/ Average investment

= 7,904/60,800

= 13%

Annual rate of return 13

%

Since average rate of return (13% ) is more than the company's required rate of return (10%), hence project is acceptable.

Accept the project.

Kindly comment if you need further assistance.

Thanks‼!